OpenStudy (purplemexican):

On the Moon, a falling object falls just 2.65 feet in the first second after being dropped. Each second it falls 5.3 feet farther than in the previous second. How far would an object fall in the first ten seconds after being dropped?

OpenStudy (welshfella):

this is a an arithmetic series where the first term is 2.65 and each subsequent term is obtained by adding the common difference 35.

OpenStudy (purplemexican):

a. 270 ft c. 313 ft b. 692 ft d. 265 ft

OpenStudy (welshfella):

nth term = a1 + ( n- 1)d where a1 = first term , n = number of terms and d = common difference so tenth term = 2.65 + 9*5.3 (common difference = 5.3 NOT 35 as i said in my first post)

OpenStudy (welshfella):

so in the second second it traveled 2.65 + 5.3 = 7.95 feet the total distance = sum of n terms = (n/2)[ a1 + L) where L = 10th term so plug in the values of a1 , n (=10) , and L and work it out

OpenStudy (purplemexican):

what the formula

OpenStudy (welshfella):

i've just given it

OpenStudy (purplemexican):

im confused please state it again if you don't mind

OpenStudy (welshfella):

plug in a1 = 2.65 , n = 10 and L = 10th term = 2.65 + 9*5.3 = 50.35

OpenStudy (purplemexican):

so L=50.35 correct?

OpenStudy (welshfella):

thats the distance travelled in the 10 th second . You want the total distance travelled after 10 seconds have passed Use the formula to work that out.

OpenStudy (welshfella):

Total distance = (n/2) * [a1 + L)

OpenStudy (welshfella):

work ouy n/2 first then a1 + L then multiply the 2 results

OpenStudy (purplemexican):

i get 265

OpenStudy (welshfella):

that's right

OpenStudy (purplemexican):

thank you

OpenStudy (welshfella):

yw